OpenStudy (snowcrystal):

Help please!!! will fan and medal if you help guide me in the right direction. Question below.

OpenStudy (snowcrystal):

What is the solution to the equation?\[3\sqrt[5]{(x+2)^{3}}+3=27\]

OpenStudy (snowcrystal):

@mathmale @math_man21

OpenStudy (snowcrystal):

What would I start with?

OpenStudy (math_man21):

for once i cant help

OpenStudy (snowcrystal):

Do you know someone who can?

OpenStudy (vhtran):

I would try working out the stuff inside the square root

OpenStudy (vhtran):

*fifth root

OpenStudy (snowcrystal):

how do i do that exactly?

OpenStudy (vhtran):

I think (you might need to FOIL but I'm not sure you can do that with a set of 3). But otherwise you can try x^3 and 2^3

OpenStudy (anonymous):

can i help.

OpenStudy (snowcrystal):

ya anyone can i just dont get this crap

OpenStudy (snowcrystal):

@pooja195

OpenStudy (snowcrystal):

should i start with the +3 thingy

OpenStudy (anonymous):

ok, \[3\sqrt[5]({x+2})^{3}=27\] \[3\sqrt[5]{(x+2)^{3}}=27-3\]

OpenStudy (snowcrystal):

okay i got that part. but i dont get the other part

OpenStudy (anonymous):

\[\sqrt[5]{(x+2)^{3}}=24/3\] \[[(x+2)^{3}]^{1/5}=8\]

OpenStudy (snowcrystal):

how did you get the 1/5

OpenStudy (anonymous):

\[(x+2)^{3}=8^{5}\] \[(x+2)=2^{3\times5/3}\]

OpenStudy (anonymous):

sqart5 means 1/5

OpenStudy (snowcrystal):

ooooooo okay

OpenStudy (anonymous):

\[(x+2)=2^{5}\] \[x+2=32\] \[x=32-2\]\[x=30\]

OpenStudy (snowcrystal):

thank you sooooooooooooo much you are a life saver!!!!!!!!!!